# Elastic displacements and viscous hydrodynamic flows in wedge-shaped geometries featuring a straight edge: Green’s functions for forces oriented parallel to the edge

### v1

@@ I Introduction
In reality, boundaries limit the system size. Solving the equations mentioned above under limiting boundaries is a significantly more challenging task. Exact analytical expressions for associated Green’s functions are available for half-space solutions that feature one flat infinitely extended surface. From there, the system is assumed to extend into the complete infinite half-space on one side of the surface. Specific surface conditions are no-slip surfaces, which do not allow for any displacement nor flow at the boundary [4, 5, 6], and free-slip conditions, which do not allow for any displacement nor flow normal to the surface but impose no restriction along the parallel direction [6]. To some extent, the situations of sheets or films of finite thickness have also been addressed, where the system is bounded by two parallel flat surfaces but still infinitely extended to the sides [7, 8, 9, 10, 11, 12, 13]. Examples of three-dimensionally bounded geometries that are genuinely finite systems and have been repeatedly studied analytically are given by spherical systems [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. The associated analytical expressions are complex. Still, the smoothness of the entire surface of the sphere favors analytical approaches. In other words, all the examples mentioned so far do not feature any kinks nor edges on their surfaces.
In the present work, we turn away from these smooth geometries and address boundaries that contain one straight infinitely extended edge. The elastic or fluid material is confined to a wedge-like geometry with the edge forming the tip of this wedge. For overdamped viscous fluid flows under conserved volume, a corresponding Green’s function has recently been derived. Yet, it was restricted to the case of free-slip surface conditions on the surfaces of the wedge and a $2\pi$-commensurate opening angle of the wedge [25]. Instead, here we generalize these considerations to linearly elastic, possibly compressible materials, coupled with no-slip and free-slip surface conditions, and allowing for arbitrary opening angles of the wedge. Our results equally apply to low-Reynolds-number flows of incompressible fluids by setting the material parameter related to compressibility in linearly elastic systems to the value that ensures incompressibility.
@@ II.1 Solution methodology
The Navier-Cauchy equation governs the displacement field $\bm{u}(\bm{x})$ of the material elements within an isotropic, homogeneous, linearly elastic medium,
@@
For $\nu=1/2$, we obtain $\sigma=1$, thereby recovering the classical form of solution proposed by Imai [26, p. 313].
The components of the displacement field in cylindrical coordinates are given by
@@
where the terms on the right-hand side arise from the coupling between components in a curvilinear coordinate system. In an infinitely extended bulk linearly elastic medium, the Green’s function reads[27]
$\widetilde{\mathcal{G}}_{ij}^{\infty}=\frac{1}{4\pi\mu\left(\sigma+1\right)}\left(\frac{\sigma}{s}\,\delta_{ij}+\frac{s_{i}s_{j}}{s^{3}}\right),$ (9)
where $s:=|\bm{s}|$ denotes the distance from the applied force singularity. Thus, for $\sigma=1$, we recover the hydrodynamic Green’s function, commonly known as the Oseen tensor [28], with $\mu$ representing the fluid dynamic viscosity [29].
@@
where $\phi_{j}^{\infty}$, with $j\in\{r,\theta,z,w\}$, are the harmonic functions in an unbounded medium, while $\phi_{j}$ represent additional fields required to satisfy additional boundary conditions.
In the following, we consider combinations of two types of boundary conditions, namely no-slip and free-slip boundary conditions. An NOS boundary condition represents a condition where there is no relative motion between the material and the boundary, that is $\bm{u}=\bm{0}$ on the boundary. An FRS boundary condition is characterized by both impermeability, where the normal component of the displacement satisfies $u_{\theta}=0$, and no-shear conditions, given by $\partial u_{r}/\partial\theta=\partial u_{z}/\partial\theta=0$.
@@ II.2 Fourier-Kontorovich-Lebedev transform
To determine the solution for the Green’s function, we use the Fourier-Kontorovich-Lebedev (FKL) transform, a technique well-suited for solving elasticity and hydrodynamics problems involving wedge-shaped geometries. This method involves transforming the axial and radial coordinates into the axial and radial wavenumbers, denoted as $k$ and ${p}$, respectively.
While the Fourier transform is widely known and its application is well-understood, the Kontorovich-Lebedev transform is much less familiar. In the late 1930s, Kontorovich and Lebedev originally introduced these transforms to solve certain boundary value problems [30, 31]. Later, further applications to boundary value problems were provided by Lebedev and Kontorovich [32], while the mathematical theory was developed by Lebedev [33, 34]. See Erdélyi et al. [35, p. 75] for further details on this transform.
@@
where $K_{i{p}}(|k|r)$ denotes the modified Bessel function of the second kind [36] of purely imaginary order $i{p}$. Note that we use the hat to denote the Fourier transform of the original function, and the tilde to represent the Kontorovich-Lebedev transform of the Fourier-transformed function. The polar angle $\theta$ remains unaffected by these transformations. Consequently, we transform the partial differential equation, Eq. (1), governing the displacement field in an elastic medium into a system of ordinary differential equations for the variable $\theta$. Note that for a positive argument, $K_{i{p}}$ yields a real number when ${p}$ is constrained to be real.
@@
for the inverse Fourier transform. In many practical situations, performing the inverse Fourier transform is much easier. Therefore, in most cases, the solution will be expressed as a single infinite integral over the radial wavenumber ${p}$, which can be evaluated using numerical methods.
@@ III.1 Free-space Green’s function
Using the parity of $\chi$ as defined by Eq. (18) with respect to $z$ in conjunction with the integral that defines the zeroth-order modified Bessel function of the second kind, see Watson [37, p. 183],
@@
We utilize the Table of Integral Transforms by Erdélyi et al. [38, p. 175] to derive the FKL transform of Eq. (21) as
@@ III.2 General form of the solution
Since $\Phi_{j}$, with $j\in\{r,\theta,z,w\}$, are harmonic functions, it follows that their corresponding FKL transforms satisfy the second-order ordinary differential equations
@@
Here, $A_{j}$ and $B_{j}$ are eight coefficients to be determined from the boundary conditions.
@@ III.3 No-slip–no-slip boundary conditions (2NOS)
imposed at $\theta=\pm\alpha$.
@@
evaluated at $\theta=\pm\alpha$. We note that $\widetilde{\phi}_{r}$ and $\widetilde{\phi}_{\theta}$ are derived from Eqs. (6), and thee derivative of $\widetilde{\phi}_{r}$ with respect to $\theta$ are given by
@@
Since $\widetilde{\phi}_{w}$ vanishes at $\theta=\pm\alpha$, it follows from the general solution given by Eq. (25) that $\widetilde{\phi}_{w}=0$ everywhere. Thus, the boundary conditions simplify to
@@
$\Delta_{\pm}=\left(\sigma\operatorname{sh}(2\alpha{p})\pm{p}\sin(2\alpha)\right)\left(\operatorname{ch}(2\alpha{p})\mp\cos(2\alpha)\right).$ (40)
@@ III.4 Free-slip–free-slip boundary conditions (2FRS)
Next, we consider the free-slip boundary condition on both surfaces to apply by imposing that
@@ III.5 No-slip–free-slip boundary condition (NOS–FRS)
Finally, we examine the case of mixed boundary conditions by imposing an NOS condition at $\theta=-\alpha$ and an FRS condition at $\theta=\alpha$. Specifically, we require Eqs. (34) to be satisfied at $\theta=-\alpha$ and Eq. (43) at $\theta=\alpha$. After simplification, we obtain for the axial components
@@ IV Green’s function in real space
$\phi_{j}=\frac{1}{\pi^{3}}\int_{-\infty}^{\infty}\mathrm{d}k\,e^{-ikz}\int_{0}^{\infty}\widetilde{\phi}_{j}\,K_{i{p}}(|k|r)\operatorname{sh}(\pi{p}){p}\,\mathrm{d}{p}\,.$ (51)
@@
see Gradshteyn and Ryzhik [41, p. 719] Eq. ET I 50(51) or the treatise by Prudnikov et al. [42, p. 390] Eq. 2.16.36 (2). Here, $P_{n}$ denotes the Legendre function of the first kind of degree $n$, with argument
@@
Similarly, we follow the same approach for the $\phi_{x}$ and $\phi_{y}$ to obtain
@@
$\mathcal{Q}_{i{p}}(r,z)=\left(\tfrac{4}{\pi}\right)^{2}\int_{0}^{\infty}\rho k\sin(kz)K_{i{p}}(k\rho)K_{i{p}}(kr)\,\mathrm{d}k,$ (58)
@@
see Gradshteyn and Ryzhik [41, p. 727] Eq. ET I 107(61) or Prudnikov et al. [42, p. 389] Eq. 2.16.36 (1). Here, $P_{n}^{m}$ denotes the associated Legendre function of the first kind of degree $n$ and order $m$.
@@
with ${}_{2}\widetilde{F}_{1}$ representing the regularized hypergeometric function, which is related to the Gaussian or ordinary hypergeometric function by the expression ${}_{2}\widetilde{F}_{1}(a,b;c;z)={}_{2}F_{1}(a,b;c;z)/\Gamma(c)$, with $\Gamma$ denoting the Euler gamma function [36]. This is the definition implemented in Maple [43]. In Mathematica [44], this is referred to as the associated Legendre function of the first kind of type 3. It differs from the standard definition in Mathematica or the one referred to as type 2, which should be multiplied by $i^{m}$ to obtain the desired result. Note that both $\mathcal{K}_{i{p}}$ and $\mathcal{Q}_{i{p}}$ yield real values.
To summarize, we have determined the harmonic functions and expressed the results as integrals over the radial wavenumber ${p}$, see Eqs. (52) and (57). The improper integrals contained in these expressions could potentially be evaluated analytically using the method of residues. However, in this study, we opt for numerical integration to evaluate them, for simplicity. The infinite integrals can be transformed into well-behaved finite integrals over the domain $[0,1]$, which are amenable to standard numerical integration methods, by introducing the change of variable $u={p}/({p}+1)$, resulting in $\mathrm{d}{p}=\mathrm{d}u/(1-u)^{2}$.
@@
Figure 2: Scaled displacement field $\bm{u}^{*}=\bm{u}/\left(F_{z}/\left(16\pi\mu\right)\right)$ in the radial-azimuthal plane at $z/\rho=1/2$ induced by a point force acting along the $z$-direction. A semi-opening angle of the wedge of $\alpha=\pi/6$ is considered, with the singularity positioned at $\beta=\alpha/2$. Here, we set $\sigma=2$. We address the different boundary conditions (a) 2NOS, (b) 2FRS, and (c) NOS–FRS.
@@
In Fig. 2, we present an exemplary scaled displacement field in the radial-azimuthal plane at $z/\rho=1/2$ for a semi-angle of the wedge of $\alpha=\pi/6$ and a singularity positioned at the polar angle $\beta=\alpha/2$. The material is assumed to be linearly elastic with $\sigma=2$. Results are shown for three types of boundary conditions (a) 2NOS, (b) 2FRS, and (c) NOS–FRS. Numerical integration is performed using $N=100$ collocation points, providing a well-resolved displacement field. In the 2FRS and NOS–FRS cases, the largest magnitudes of displacement are observed near polar azimuthal coordinates of the singularity, $(r,\theta)\simeq(\rho,\beta)$. Contrarily, for the 2NOS case, the region of larger displacement is shifted toward $\theta\simeq 0$. Apparently, the no-slip surface condition enforced on both boundaries shifts the location of the maximum displacement away from the boundary of the wedge towards the center.
Figure 3: Displacement field in the radial-axial plane of $\theta=0$ induced by a point force acting along the $z$-direction. The parameters are identical to those in Fig. 2. Again, the scaled displacement field is defined as $\bm{u}^{*}=\bm{u}/\left(F_{z}/\left(16\pi\mu\right)\right)$. We examine the different boundary conditions (a) 2NOS, (b) 2FRS, and (c) NOS–FRS.
@@
Figure 3 presents the corresponding results in the radial-axial plane using the same set of parameters as in Fig. 2. Here, the plots are shown in the azimuthal plane of $\theta=0$. We observe that the pattern induced by a point force exhibits similarities between 2NOS and NOS–FRS conditions. However, the overall magnitude of the displacement field for the 2NOS condition is lower than under NOS–FRS conditions. Illustratively, no-slip conditions are stricter and additionally constrain the displacement field.

### v2

# Elastic displacements and viscous hydrodynamic flows in wedge-shaped geometries with a straight edge: Green’s functions for parallel forces

@@ I Introduction
In reality, boundaries limit the system size. Solving the equations mentioned above under limiting boundaries is a significantly more challenging task. Exact analytical expressions for associated Green’s functions are available for half-space solutions that feature one flat, infinitely extended surface. From there, the system is assumed to extend into the complete infinite half-space on one side of the surface. Specific surface conditions are no-slip surfaces, which do not allow for any displacement nor flow at the boundary [4, 5, 6], and free-slip conditions, which do not allow for any displacement nor flow normal to the surface but impose no restriction along the parallel direction [6]. To some extent, the situations of sheets or films of finite thickness have also been addressed, where the system is bounded by two parallel flat surfaces but still infinitely extended to the sides [7, 8, 9, 10, 11, 12, 13]. Examples of three-dimensionally bounded geometries that are genuinely finite systems and have been repeatedly studied analytically are given by spherical systems [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. The associated analytical expressions are complex. Still, the smoothness of the entire surface of the sphere favors analytical approaches. In other words, all the examples mentioned so far do not feature any kinks nor edges on their surfaces.
In the present work, we turn away from these smooth geometries and address boundaries that contain one straight infinitely extended edge. The elastic or fluid material is confined to a wedge-like geometry with the edge forming the tip of this wedge. Flows at low Reynolds numbers in incompressible viscous fluids near three-dimensional corners were originally investigated by Sano and Hasimoto, who examined a series of progressively intricate wedge-shaped geometries [25, 26, 27, 28, 29]. More recently, the three-dimensional Stokes flow near a corner has been reexamined using techniques from complex analysis [30, 31]. For overdamped viscous fluid flows under conserved volume, a corresponding Green’s function has recently been derived. Yet, it was restricted to the case of free-slip surface conditions on the surfaces of the wedge and a $2\pi$-commensurate opening angle of the wedge [32]. Instead, here we generalize these considerations to linearly elastic, possibly compressible materials, coupled with no-slip and free-slip surface conditions, and allowing for arbitrary opening angles of the wedge. Our results equally apply to low-Reynolds-number flows of incompressible fluids by setting the material parameter related to compressibility in linearly elastic systems to the value that ensures incompressibility. Together with the solution obtained in Ref. 33 for the case of a perpendicular point force, this allows the construction of the general solution for an arbitrary force direction.
@@ II.1 Solution methodology
The Navier-Cauchy equation governs the displacement field $\bm{u}(\bm{x})$ of the material elements within an isotropic, homogeneous, linearly elastic medium. All transients of applying a constant force density are assumed to have decayed, so that the resulting static deformed state is described by
@@
For $\nu=1/2$, we obtain $\sigma=1$, thereby recovering the classical form of solution proposed by Imai [34, p. 313].
The problem can be solved more conveniently using a cylindrical coordinate system by expressing the displacement field in terms of its radial, azimuthal, and axial components. They are denoted as $u_{r}$, $u_{\theta}$, and $u_{z}$, respectively. Applying the standard relations of coordinate transformation from Cartesian to cylindrical coordinates, the components of the displacement field in the cylindrical system are given by
@@
where the terms on the right-hand side arise from the coupling between components in a curvilinear coordinate system. In an infinitely extended bulk linearly elastic medium, the Green’s function corresponding to a point force applied at position $\bm{x}_{0}$ reads[35]
$\mathcal{G}_{ij}^{\infty}=\frac{1}{4\pi\mu\left(\sigma+1\right)}\left(\frac{\sigma}{s}\,\delta_{ij}+\frac{s_{i}s_{j}}{s^{3}}\right),$ (9)
where $\bm{s}=\bm{x}-\bm{x}_{0}$ and $s:=|\bm{s}|$ denotes the distance from the applied force singularity. Thus, for $\sigma=1$, we recover the hydrodynamic Green’s function, commonly known as the Oseen tensor [36], with $\mu$ representing the shear viscosity [37].
@@
where $\phi_{j}^{\infty}$, with $j\in\{r,\theta,z,w\}$, are the harmonic functions in an unbounded medium. The fields $\phi_{j}$ represent unknown functions to be determined, subject to the boundary conditions imposed at the boundaries of the wedge.
In the following, we consider combinations of two types of boundary conditions, namely no-slip and free-slip boundary conditions. A NOS boundary condition represents a condition where there is no relative motion between the material and the boundary, that is $\bm{u}=\bm{0}$ on the boundary. A FRS boundary condition is characterized by both impermeability, where the normal component of the displacement satisfies $u_{\theta}=0$, and no-shear conditions, given by $\partial u_{r}/\partial\theta=\partial u_{z}/\partial\theta=0$.
@@ II.2 Fourier-Kontorovich-Lebedev transform
To determine the solution for the Green’s function, we use the Fourier-Kontorovich-Lebedev (FKL) transform, a technique well-suited for solving elasticity and hydrodynamics problems involving wedge-shaped geometries. This method involves transforming the axial and radial coordinates into the axial and radial wavenumbers, denoted as $k$ and p{p}, respectively.
On one hand, the Fourier transform is typically applied only with respect to the $z$-coordinate. Frequently, corresponding systems are unbounded in the $z$-direction and exhibit translational invariance, making it natural to use the Fourier transform to handle these infinite or semi-infinite domains. Consequently, differential operators in $z$ are transformed into algebraic ones, which simplifies the analysis. On the other hand, the radial coordinate $r$ ranges from zero to infinity, which makes the Kontorovich–Lebedev (KL) transform well-suited for this variable. The KL transform is effective for radial problems involving wedge or conical geometries, because it diagonalizes operators that involve Bessel functions, which naturally arise in cylindrical coordinates. Combining the two transforms allows for a separation of variables that aligns with the considered geometry and the associated boundary conditions. The Fourier transform reduces the dimensionality by handling the axial direction. Simultaneously, the KL transform efficiently manages the radial dependence in a manner consistent with the wedge boundaries and the spectral properties of the Laplacian in the radial coordinate. This combined approach leverages the strengths of each transform for the respective variable.
While the Fourier transform is widely known and its application is well-understood, the KL transform is much less familiar. In the late 1930s, Kontorovich and Lebedev originally introduced these transforms to solve certain boundary value problems [38, 39]. Later, further applications to boundary value problems were provided by Lebedev and Kontorovich [40], while the mathematical theory was developed by Lebedev [41, 42], see Erdélyi et al. [43, p. 75] for further details on this transform. The KL transform has been extensively applied to a range of physical problems involving wedge geometries, particularly in the context of electromagnetic scattering and diffraction [44, 45, 46, 47, 48, 49, 50, 51, 52], as well as in the study of fluid flows [53, 54]. Other studies have employed the KL transform in the presence of mixed boundary conditions [55, 56].
@@
where $K_{i{p}}(|k|r)$ denotes the modified Bessel function of the second kind [57] of purely imaginary order ipi{p}. Note that we use the hat to denote the Fourier transform of the original function, and the tilde to represent the Kontorovich-Lebedev transform of the Fourier-transformed function. The polar angle $\theta$ remains unaffected by these transformations. Consequently, we transform the partial differential equation, Eq. (1), governing the displacement field in an elastic medium into a system of ordinary differential equations for the variable $\theta$. Note that for a positive argument, $K_{i{p}}$ yields a real number when p{p} is constrained to be real.
@@
for the inverse Fourier transform. In many practical situations, performing the inverse Fourier transform is much easier. Therefore, in most cases, the solution will be expressed as a single infinite integral over the radial wavenumber p{p}, which can be evaluated using numerical methods.
@@ III.1 Free-space Green’s function
Using the parity of $\chi$ as defined by Eq. (18) with respect to $z$ in conjunction with the integral that defines the zeroth-order modified Bessel function of the second kind, see Watson [58, p. 183],
@@
We utilize the Table of Integral Transforms by Erdélyi et al. [59, p. 175] to derive the FKL transform of Eq. (21) as
@@ III.2 General form of the solution
Since $\Phi_{j}$, with $j\in\{x,y,z,w\}$, are harmonic functions, it follows that their corresponding FKL transforms satisfy the second-order ordinary differential equations
@@
Here, $A_{j}$ and $B_{j}$ are eight coefficients to be determined from the boundary conditions. It is important to note that $\widetilde{\phi}_{r}$ and $\widetilde{\phi}_{\theta}$ are not harmonic functions, see Eq. (8). Thus, to them the form of solutions in Eq. (25) does not apply.
@@ III.3 No-slip–no-slip boundary conditions (2NOS)
imposed at $\theta=\pm\alpha$. We could have chosen alternative representations, but given the flexibility in selecting harmonic functions as discussed above, we opted for the form of solution that we consider simplest.
@@
evaluated at $\theta=\pm\alpha$. We note that $\widetilde{\phi}_{r}$ and $\widetilde{\phi}_{\theta}$ are derived from Eqs. (6), and the derivative of $\widetilde{\phi}_{r}$ with respect to $\theta$ is given by
@@
Since $\widetilde{\phi}_{w}$ vanishes at $\theta=\pm\alpha$, it follows from the general solution given by Eq. (25) that $\widetilde{\phi}_{w}=0$ everywhere. However, this is not the case for $\widetilde{\phi}_{r}$, because Eq. (25) does not apply to $j=r$. Thus, the boundary conditions simplify to
@@
$\hskip-2.84544pt\Delta_{\pm}=\left(\sigma\operatorname{sh}(2\alpha{p})\pm{p}\sin(2\alpha)\right)\left(\operatorname{ch}(2\alpha{p})\mp\cos(2\alpha)\right).$ (40)
@@ III.4 Free-slip–free-slip boundary conditions (2FRS)
Next, we impose the free-slip boundary condition on both surfaces by requiring that
@@ III.5 No-slip–free-slip boundary condition (NOS–FRS)
Finally, we examine the case of mixed boundary conditions by imposing a NOS condition at $\theta=-\alpha$ and a FRS condition at $\theta=\alpha$. Specifically, we require Eqs. (34) to be satisfied at $\theta=-\alpha$ and Eq. (43) at $\theta=\alpha$. After simplification, we obtain for the axial components
@@ IV Green’s function in real space
$\hskip-2.84544pt\phi_{j}=\frac{1}{\pi^{3}}\int_{-\infty}^{\infty}\mathrm{d}k\,e^{-ikz}\int_{0}^{\infty}\widetilde{\phi}_{j}\,K_{i{p}}(|k|r)\operatorname{sh}(\pi{p}){p}\,\mathrm{d}{p}\,.$ (51)
@@
see Gradshteyn and Ryzhik [60, p. 719] Eq. ET I 50(51) or the treatise by Prudnikov et al. [61, p. 390] Eq. 2.16.36 (2). Here, $P_{n}$ denotes the Legendre function of the first kind of degree $n$, with argument
@@
Similarly, we follow the same approach for $\phi_{x}$ and $\phi_{y}$ to obtain
@@
$\hskip-2.84544pt\mathcal{Q}_{i{p}}(r,z)=\left(\tfrac{4}{\pi}\right)^{2}\int_{0}^{\infty}\rho k\sin(kz)K_{i{p}}(k\rho)K_{i{p}}(kr)\,\mathrm{d}k,$ (58)
@@
see Gradshteyn and Ryzhik [60, p. 727] Eq. ET I 107(61) or Prudnikov et al. [61, p. 389] Eq. 2.16.36 (1). Here, $P_{n}^{m}$ denotes the associated Legendre function of the first kind of degree $n$ and order $m$.
@@
with $_{2}\widetilde{F}_{1}$ representing the regularized hypergeometric function, which is related to the Gaussian or ordinary hypergeometric function by the expression $_{2}\widetilde{F}_{1}(a,b;c;z)={}_{2}F_{1}(a,b;c;z)/\Gamma(c)$, with $\Gamma$ denoting the Euler gamma function [57]. This is the definition implemented in Maple [62]. In Mathematica [63], this is referred to as the associated Legendre function of the first kind of type 3. It differs from the standard definition in Mathematica or the one referred to as type 2, which should be multiplied by $i^{m}$ to obtain the desired result. We note that both $\mathcal{K}_{i{p}}$ and $\mathcal{Q}_{i{p}}$ yield real values. Far from the singularity, $\mathcal{K}_{ip}$ and $\mathcal{Q}_{ip}$, along with their derivatives, vanish, thereby ensuring that the solution satisfies the regularity conditions at infinity.
Figure 2: Scaled displacement field $\bm{u}^{*}=\bm{u}/\left(F_{z}/\left(16\pi\mu\rho\right)\right)$ in the radial-azimuthal plane at $z/\rho=1/2$ induced by a point force acting along the $z$-direction. A semi-opening angle of the wedge of $\alpha=\pi/6$ is considered, with the singularity positioned at $\beta=\alpha/2$. Panels shown in the top row (a)–(c) correspond to $\sigma=1$, and those in the bottom row (d)–(f) to $\sigma=2$. We address the different boundary conditions (a,d) 2NOS, (b,e) 2FRS, and (c,f) NOS–FRS.
To summarize, we have determined the harmonic functions and expressed the results as integrals over the radial wavenumber p{p}, see Eqs. (52) and (57). The improper integrals contained in these expressions could potentially be evaluated analytically using the method of residues. However, in this study, we opt for numerical integration to evaluate them, for simplicity. The infinite integrals can be transformed into well-behaved finite integrals over the domain $[0,1]$, which are amenable to standard numerical integration methods, by introducing the change of variable u=p/(p+1)u={p}/({p}+1), resulting in $\mathrm{d}{p}=\mathrm{d}u/(1-u)^{2}$.
@@
Figure 3: Displacement field in the radial-axial plane of $\theta=0$ induced by a point force acting along the $z$-direction. The parameters are identical to those in Fig. 2. Again, the scaled displacement field is defined as $\bm{u}^{*}=\bm{u}/\left(F_{z}/\left(16\pi\mu\rho\right)\right)$. The top-row panels (a)–(c) correspond to $\sigma=1$, while the bottom-row panels (d)–(f) correspond to $\sigma=2$. Considered boundary conditions are (a,d) 2NOS, (b,c) 2FRS, and (c,f) NOS–FRS.
@@
Figure 4: Displacement field (a)–(c) in the radial–azimuthal plane at $z/\rho=1/2$, and (d)–(f) in the radial–axial plane at $\theta=0$, all for $\sigma=1$. The displacements are induced by a point force acting in the $z$-direction. Remaining parameters are identical to those used in Fig. 2 for $\sigma=1$. Here, the depicted scaled displacement field is defined as $\bm{U}^{*}=(\bm{u}-\bm{u}^{\infty})/(F_{z}/(16\pi\mu\rho))$. It therefore shows the contributions to the displacement field solely induced by the presence of the boundaries, that is, we subtract the solution for a three-dimensionally infinitely extended domain. Considered boundary conditions are (a,d) 2NOS, (b,e) 2FRS, and (c,f) NOS–FRS.
@@
In Fig. 2, we present an exemplary scaled displacement field in the radial-azimuthal plane at $z/\rho=1/2$ for a semi-angle of the wedge of $\alpha=\pi/6$ and a singularity positioned at the polar angle $\beta=\alpha/2$. The material is assumed to be linearly elastic with $\sigma=1$ (incompressible) for the top row and $\sigma=2$ $(\nu=1/4)$ for the bottom row. Results are shown for three types of boundary conditions: 2NOS in (a) and (d), 2FRS in (b) and (e), and NOS–FRS in (c) and (f). Numerical integration is performed using $N=100$ collocation points, providing a well-resolved displacement field. In the 2FRS and NOS–FRS cases, the largest magnitudes of displacement are observed near polar azimuthal coordinates of the singularity, $(r,\theta)\simeq(\rho,\beta)$. Contrarily, for the 2NOS case, the region of larger displacement is shifted toward $\theta\simeq 0$. Apparently, the no-slip surface condition enforced on both boundaries shifts the location of the maximum displacement away from the boundary of the wedge towards the center. Overall, the case $\sigma=1$ results in magnitudes of displacement approximately twice as large as those for $\sigma=2$. For incompressible materials ($\sigma=1$), imposed displacements cannot be accommodated by local compressions or dilations. Given that $\sigma=1$ also formally corresponds to the situation of low-Reynolds-number hydrodynamics of incompressible fluids, these results are particularly useful for validating numerical methods. Specific examples include those based on the boundary element method with a meshed wedge-like boundary [64].
@@
Figure 3 presents the corresponding results in the radial-axial plane using the same set of parameters as in Fig. 2. Here, the plots are shown in the azimuthal plane of $\theta=0$. We observe that the pattern induced by a point force exhibits similarities between 2NOS and NOS–FRS conditions. However, the overall magnitude of the displacement field for the 2NOS condition is lower than under NOS–FRS conditions. Illustratively, no-slip conditions are stricter and additionally constrain the displacement field. In the 2NOS case, pronounced vortex structures are observed, particularly for $\sigma=1$. These vortices represent closed lines of displacement that arise due to the boundary conditions, significantly influencing the local appearance. Remarkably, the displacement field in this case exhibits two distinct maxima symmetrically located about the plane $z=0$. As $\sigma$ is increased to $2$, these maxima shift toward $z=0$ and merge.
In Fig. 4, we present that part of the solution required to satisfy the imposed boundary conditions, using the same parameters as in Fig. 2 for $\sigma=1$. That is, we subtract from the solution $\bm{u}$ the displacement field $\bm{u}^{\infty}$ for a three-dimensionally infinitely extended system. The results are shown (a)–(c) in scaled form in the radial–azimuthal plane at $z/\rho=1/2$, and (d)–(f) in the radial–axial plane at $\theta=0$. In the radial–azimuthal plane in (a)–(c), the displacement field is oriented away from the boundaries of the wedge. When combined with the infinite-space solution in an unbounded elastic medium, this ensures that the overall field satisfies the boundary conditions. In the radial–axial plane in (d)–(f), the displacement field descends for 2NOS and NOS–FRS, while for 2FRS it ascends, following the direction of the applied point force. This behavior results from enforcing no-slip boundary conditions at one or both boundaries. We mention that in the context of low-Reynolds-number hydrodynamics ($\sigma=1$) it is well known that boundary-induced disturbances of the flow can cause migration toward or away from boundaries for polymers [65], non-spherical particles [66], or self-propelled active particles and microswimmers [67, 68, 69].